An artificial lake is in the shape of a rectangle and has an area of 9/20 square mile. The width of the lake is 1/5 the length of the lake. What are the dimensions of the lake?
Answer by stanbon(75887)

An artificial lake is in the shape of a rectangle and has an area of 9/20 square mile. The width of the lake is 1/5 the length of the lake. What are the dimensions of the lake?
----------------------
Let the length be "L"
Then the width = (1/5)L
--------------------------
Equation:
L(L/5) = 9/20
-----
L^2 = 9/4
-----
Lenght = 3/2 = 1.5 miles
Width = (1/5)(3/2) = 3/10 = 0.3 miles
-----------------------------------------
Cheers,
Stan H.
-------------------